OpenStudy (anonymous):

the curve y=ax2+bx+c passes through the point(1,2) and is tangent to the line y=x at the origin.find a,b and c..........in ax2......2 is a power of x

OpenStudy (anonymous):

Eh, here's something you can do. You can put in the given information like this: (2)=a(1)^2+b(1)+c (0)=a(0)^2+b(0)+c Then find a, b, and c such that ax^2+bx+c=x has only one solution (can use discriminant test or something like that).

OpenStudy (anonymous):

no...we have to solve it by taking derivative rule and slope.........but how????

OpenStudy (anonymous):

Derivative of quadratic gives slope of tangent line

OpenStudy (anonymous):

Oh, if you know calculus, that'll be a lot easier.

OpenStudy (anonymous):

and the y intercept for the quadratic is pretty clear

OpenStudy (anonymous):

what @estudier is telling you is if the line \(y=x\) is tangent at the origin, then the value of function at the origin must be the same a \(y=x\) at the origin, namely \(0\) so you know \(c=0\)

OpenStudy (anonymous):

Take derivative of ax^2+bx. You know c has to =0, so just leave it out.

OpenStudy (anonymous):

could any solve it and tell me the answer of a and b????

OpenStudy (anonymous):

You can do it. What is the first derivative of ax^2+bx?

OpenStudy (anonymous):

2ax+b

OpenStudy (anonymous):

Derivative equals slope of the tangent line, so set that equal to the slope of the tangent line.

OpenStudy (anonymous):

we would use the value of x in 2ax that is one....right????

OpenStudy (anonymous):

The value of x at the point of tangency is 0. The slope of the tangent line is 1.

OpenStudy (anonymous):

no...i am talking about the curve that passes through the point (1,2)

OpenStudy (anonymous):

Yes, me too.

OpenStudy (anonymous):

"...and is tangent to the line y=x at the origin." The line is tangent at x=0.

OpenStudy (anonymous):

so i think b=-1...

OpenStudy (anonymous):

Derivative = slope Derivative = 2ax+b Slope = 1 2ax+b=1 x=0 at point of tangency. 2a(0)+b=1 Yes?

OpenStudy (anonymous):

yeah....but what about a????

OpenStudy (anonymous):

Now you can use the point (1,2) to find that.

OpenStudy (anonymous):

oh yes.....its done a=1.....thankyou.....

OpenStudy (anonymous):

You're welcome.